Find the standard matrices A and A for T = T2 o T1 and T = T1 o T2. T1: R2 → R2, T1(x, y) = (x – 4y, 3x + 3y) T2: R2 → R2,
Answer #1

sull given ban gaan To Cosy) = (2-49, 34y) x-49, Transformation metrix corresponding to Ti TE 1 3 -4 3 similary to x, y) = (13x+3y transition metrix corresponding to A zSinul aily. A (27) = Ti (12 XX, Y)) TO A (.7) = transition Métux corresponding to A (ay) Lo 3.1